Algebra- solving sytems of equations in three varaiables:
( Answer is needed within 1.5 hours, by 11"30 AM Pacific time - Thanks!)
3x+5y+4z =13
5z+2y+3z=-9
6x+3y+4z=-8

Answer by CharlesG2(834)

Algebra- solving sytems of equations in three varaiables:
( Answer is needed within 1.5 hours, by 11"30 AM Pacific time - Thanks!)
3x+5y+4z =13
5z+2y+3z=-9 (--> assuming this is 5x + 2y + 3z = -9)
6x+3y+4z=-8

using method of elimination, choosing z first since I see two 4z's

use 1st equation and multiply 3rd equation by -1

3x + 5y + 4z = 13
-6x - 3y - 4z = 8
add these to get -3x + 2y = 21

use 1st equation multiplied by 3 and 2nd equation multipled by -4

9x + 15y + 12z = 39
-20x - 8y - 12z = 36
add these to get -11x + 7y = 75

we now have:
-3x + 2y = 21 (multiply this one by 7)
-11x + 7y = 75 (multiply this one by -2)

to get:
-21x + 14y = 147
22x - 14y = -150
add these to get x = -3

plug x = -3 into -3x + 2y = 21
-3(-3) + 2y = 21
9 + 2y = 21
2y = 12
y = 6

now going back to the original equations, plug x = -3 and y = 6
into 3x + 5y + 4z = 13
3(-3) + 5(6) + 4z = 13
-9 + 30 + 4z = 13
21 + 4z = 13
4z = -8
z = -2

solution is (-3,6,-2)

hope this helps you